OpenStudy (anonymous):

7x over 8 = 21 24 18.375 1.857 19

OpenStudy (anonymous):

21 times 8 over 7.

OpenStudy (anonymous):

\[\frac{7x}{8}=21\\8*\frac{7x}{8}=8*21 \\7x=168\\ \frac{7x}{7}= \frac{168}{7} \]

OpenStudy (anonymous):

so my answer would be 24

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thank You